OpenStudy (anonymous):

Sasha solved an equation, as shown below: Step 1: 8x = 56 Step 2: x = 56 – 8 Step 3: x = 48 Is Sasha's solution correct or incorrect? If the solution is incorrect, explain why it is incorrect and show the correct steps to solve the equation.

OpenStudy (anonymous):

@icecreamman

OpenStudy (anonymous):

hi payton! so first of all how would you solve this problem?

OpenStudy (anonymous):

um okay, i feel stupid. that is the wrong question. heres the real one; 5(2x – 3) = 5 Part A: How many solutions does this equation have? Part B: What are the solutions to this equation? Show your work. i need help showing my work

OpenStudy (anonymous):

ok so do you know how to solve this problem?

OpenStudy (anonymous):

or start to solve this problem?

OpenStudy (anonymous):

simplify it? im not so good at math.

OpenStudy (anonymous):

i need alot of help

OpenStudy (anonymous):

ok so the first step is to distribute. 5(2x - 3) = 5 so multiply everything out of the parenthesis. do you know how to do that?

OpenStudy (anonymous):

@paytoncobain

OpenStudy (anonymous):

5 times 2x and 5 times 3?

OpenStudy (anonymous):

5*2x and 5*-3 right? so what does that equal to?

OpenStudy (anonymous):

@paytoncobain

OpenStudy (anonymous):

10x and -15

OpenStudy (anonymous):

@paytoncobain

OpenStudy (anonymous):

yes so now the equation would be 10x - 15 = 5 right?

OpenStudy (anonymous):

okay, and now what?

OpenStudy (anonymous):

now you have to comvine like terms. do you know how to do that?

OpenStudy (anonymous):

-15 and 5 right?

OpenStudy (anonymous):

yes! so combine that and what do you get?

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

right but can you give me the whole equation

OpenStudy (anonymous):

10x-10=5?

OpenStudy (anonymous):

@icecreamman